# Active Learning-Based Optimization of Hydroelectric Turbine Startup to Minimize Fatigue Damage

### v1

@@ 1 Introduction
Among these conditions, HGU startup sequences induce substantial stress on the runner (Gagnon et al., 2010). This poses a significant challenge for hydropower utilities, particularly in light of the anticipated increase in starting cycle frequency (Savin et al., 2020). Over the lifetime of the turbine, the stress cycles experienced by the runner contribute to fatigue, leading to crack propagation (Gummer and Etter, 2008). The risk of blade cracking increases the frequency of the required inspections, which results in significant financial costs, reduced flexibility, and a loss of power output for hydropower utilities (Nilsson and Sjelvgren, 1997). Therefore, mitigating turbine stresses during startup is crucial to enhance the reliability and longevity of hydroelectric plants in the evolving energy landscape as shown in Gagnon et al. (2020).
@@
Stress levels experienced by hydroturbines during startup are closely linked to the trajectory in terms of guide vane opening and rotational speed (Čepa et al., 2018; Mukai et al., 2022). The guide vane opening sequence is parameterized and ensures that the turbine reaches synchronous speed while respecting specific operational requirements. Optimization of startup sequence parametrization can be approached through experimental reduced-scale model tests (Seydoux et al., 2024; Schmid et al., 2022) and on-site campaigns (Unterluggauer et al., 2020; Čepa et al., 2018). Model testing allows for cost reductions and exploration of a wide variety of startup scenarios without the access and time constraints encountered in prototype tests. However, the transposition of turbine mechanical responses during startup from model to full-scale remains a significant challenge, and no rigorous methodology has been established to date. In contrast, prototype measurements provide insights into the real mechanical behavior of full-scale turbines. However, these measurements are challenging and costly for hydroelectric operators due to time constraints and the harsh environment in which hydroelectric turbines operate, which often leads to the rapid degradation of strain gauges. Therefore, startup optimization must rely on the limited data acquired during short, dedicated measurement campaigns.
@@
Gagnon et al. (2014, 2016) presented the results of such campaigns, highlighting the effects of startup sequences on blade stresses. Based on their results, they defined optimal startup scenarios in terms of stress levels. However, these scenarios are mainly derived through trial-and-error processes and expert experience, leaving the potential for further optimization untapped. While optimization may be performed post-campaign using a virtual sensor approach (Gagnon et al., 2023), the results cannot be validated without a subsequent experimental campaign.
@@
To address these limitations, we propose a novel methodology that allows for the optimization of startup trajectory parameters during a measurement campaign, ultimately contributing to extending the life of hydroturbines. By combining an active learning framework with virtual sensors and black-box optimization, the active learning algorithm sequentially generates various startup parameters to test on the turbine. At each step, new data are used to refine the model and optimization process. This approach was successfully tested during an experimental campaign conducted on a full-scale Francis turbine instrumented with several strain gauge rosettes.
@@
This article is structured as follows. In Section 2, the problem is formulated in details, from HGU dynamics and strain generation to the resulting active optimization problem. Section 3 presents the proposed methodology, which combines an HGU dynamics simulator, virtual strain gauges, black-box optimization and the active learning approach. This method was applied to a real HGU during a measurement campaign, and the results are provided and discussed in Section 4.
@@ 2 Problem description
At time $t$, the strain $s(t)$ at a given location of a turbine blade can be measured by strain gauges, usually in microstrains ($\mu$S), at frequency $f_{M}$. During a startup process, the measured trajectory can be captured as $\tau_{M}^{i}$, defined as the combination of the dynamic trajectory $\tau_{D}^{i}$ and the strain measurements $s^{i}_{n}$:
@@
On turbine blades, repeated high-amplitude strain cycles lead to fatigue and potential cracks (Gagnon and Thibault, 2015). The greater the cycle amplitude, the greater the accumulated damage to the blade over time. Therefore, in fatigue analysis, the largest amplitude strain/stress cycle, meaning the difference between the highest and lowest strain values over the complete start process, has the most impact on fatigue. As a result, the objective of minimizing fatigue during startup can be approximated by minimizing the amplitude of the largest strain cycle. Formally, this can be expressed as minimizing the following loss over a given trajectory:
$\mathcal{L}(\tau_{M}^{i})=\max_{n\in\nu_{M}^{i}}(s_{n})-\min_{n\in\nu_{M}^{i}}(s_{n})$ (4)
@@ 3.1 Startup trajectory simulation
Turbine dynamics can be addressed using the SIMSEN software (Nicolet, 2007). A SIMSEN model of the HGU includes the upstream and downstream reservoirs, the water piping system, and the turbine. It can be used to simulate the rotational speed $\omega(t)$ for a predefined guide vane opening $o(t)$ time series. A typical simulation of a single HGU startup takes between 30 seconds and 1 minute. This is problematic for an active optimization framework, where hundreds of simulations must be performed in a limited time. Moreover, in practice, the guide vane opening $o(t)$ is not known beforehand; instead, it is calculated as part of the speed governor model, using the rotational speed $\omega(t)$. To properly address this bidirectional coupling, a co-simulation approach would have to be used in order to ensure communication between the SIMSEN model and speed governor model at each time step, further increasing computational time.
@@ 3.3 Strain modeling with virtual sensors
Evaluating the objective in Equation (16), requires an estimation of $b^{u}$ and $b^{l}$. Given strain measurement data, it is possible to build a model of these distributions using the virtual sensor approach (Gagnon et al., 2023). $b^{u}(\cdot|\omega_{n},o_{n})$ and $b^{l}(\cdot|\omega_{n},o_{n})$ are assumed to be conditional normal distributions, parameterized by mean $\mu^{u}_{n}$ and $\sigma^{u}_{n}$ (respectively, $\mu^{l}_{n}$ and $\sigma^{l}_{n}$). These parameters are estimated using the deep learning model $m_{\phi}(\omega_{n},o_{n})=\left(\hat{\mu}^{u}_{n},\hat{\sigma}^{u}_{n},\hat{\mu}^{l}_{n},\hat{\sigma}^{l}_{n}\right)$ trained on a dataset $\mathcal{D}$ of enveloped trajectories $\{\tau_{e}\}$.
Model $m_{\phi}$ is parameterized as a feed-forward neural network composed of one input layer of dimension 2 ($\omega_{n},o_{n})$, two hidden layers of dimension 32, and one output layer of dimension 4 (two $\mu$ values and two $\sigma$ values). The hidden layers are normalized using batch normalization (Ioffe and Szegedy, 2015). Because the trajectories are time series, it could seem intuitive to use recurrent neural networks such as LSTM or GRU and condition the prediction over the past trajectory $\tau_{D}$. However, such architectures do not produce better results experimentally, hinting that, during startups, the current valve opening and rotational speed values contain the necessary information to predict the strain envelope.
@@
Since $b^{u}$ and $b^{l}$ are modeled as conditional normal distributions, $m_{\phi}$ was trained using the $\beta$-negative log-likelihood loss function (Seitzer et al., 2022), which was shown to robustly estimate heteroscedastic aleatoric uncertainty. Note that predicting $\mu^{u}_{n}$ and $\sigma^{u}_{n}$ is equivalent to estimating the aleatoric uncertainty of $s^{u}_{n}$ (and respectively for $s^{l}_{n}$). Model $m_{\phi}$ was trained over three epochs, using the Adam optimizer with a learning rate of 0.001 and a mini-batch size of 32.
@@ 3.4 Black-box optimization
Given a virtual sensor $m_{\phi}(\omega_{n},o_{n})$, the objective of Equation (16) under the time constraint of Equation (6) can be addressed using black-box optimization. We use the NOMAD 4 (Audet et al., 2022) with a budget of $N_{I}$ calls to the black-box.
@@
$c_{c}(\theta^{i})=\begin{cases}0&\text{if }t^{i}_{\mathrm{st}}<0.5T_{\mathrm{st}}\\ 0.05(t^{i}_{\mathrm{st}}-0.5T_{\mathrm{st}})/(0.5T_{\mathrm{st}})&\text{if }0.5T_{\mathrm{st}}\leq t^{i}_{\mathrm{st}}<T_{\mathrm{st}}\\ 1+\left(t^{i}_{\mathrm{st}}-T_{\mathrm{st}}\right)/(0.2T_{\mathrm{st}})&\text{if }t^{i}_{\mathrm{st}}\geq T_{\mathrm{st}}\\ \end{cases}$ (19)
In the second case, the time constraint is respected. However, $c_{c}$ signals to NOMAD that, given equal strain amplitude, a shorter trajectory is preferred. The last case, representing a violation of the time constraint, is set as an indicator of 1, plus a penalty that increases as the scale of the constraint violation increases to guide NOMAD towards shorter trajectories. If $\omega$ has not stabilized at $\omega_{S}$ after $2T_{\mathrm{st}}$, the simulation is stopped and $t^{i}_{\mathrm{st}}$ is set to $2T_{\mathrm{st}}$.
To ensure compliance with the time constraint, $c_{c}(\theta^{i})$ in the event of a time constraint violation must be higher than any possible strain-related cost $\alpha_{d}c_{s}(\theta^{i})$. Therefore, $\alpha_{d}$ is set as the inverse of the difference between the maximum and minimum strain values encountered in the whole dataset $\mathcal{D}$. This ensures that $\alpha_{d}c_{s}(\theta^{i})\leq 1$ in all cases while $c_{c}(\theta^{i})\geq 1$ if $t^{i}_{\mathrm{st}}\geq T_{\mathrm{st}}$.
@@ 3.5 Active learning and optimization
The proposed active scheme is presented in Figure 6, where the black box optimizer is used at every outer loop iteration to provide the parameters $\theta^{j+1}$ to test on the turbine. The new measured trajectory is then added to the dataset, and the virtual sensor is retrained with the new data. The process goes through three phases with predetermined trajectory budgets: initialization ($N_{init}$), active learning ($N_{act}$), and optimization ($N_{opt}$), such that $N_{init}+N_{act}+N_{opt}=N_{\mathrm{b}}$. For the remainder of the paper, the variables indexed by $i$ in the single black-box loop will be indexed by ${i,j}$ to include nested loop index $j$.
@@
The objective of the active learning phase is to explore the space of trajectories in order to supply the most relevant data for training the virtual sensor. To do so, the same black-box optimization system as presented in Section 3.4 is used. However, the evaluator is modified to account for the epistemic uncertainty, expressed as the standard deviation $\sigma_{\mathrm{ep},n}^{\{u,l\},i,j}$ of the ensemble’s individual $m$ model predictions of sum $\hat{\mu}^{\{u,l\},i,j}_{n}+\hat{\sigma}^{\{u,l\},i,j}_{n}$. Inspired by the upper confidence bound approach (Auer et al., 2002) for exploration, the strain cost in Equation (18) is modified as follows:
@@ 4.1 Experimental setup
Our case study is an instrumented Francis turbine runner installed in a storage hydroelectric power plant. The specific speed of the turbine is $n_{QE}=nQ^{0.5}/(gH)^{3/4}=65$, with $n$, the runner frequency, $Q$, the nominal flow rate, $g$, the gravity and $H=72$ m, the nominal head. For this power plant, the time limit for reaching synchronous speed was set to $T_{st}=90$s.
Two blades were instrumented with four strain gauge rosettes near the trailing edge, with two rosettes positioned on the suction side and two on the pressure side. An additional uniaxial strain gauge was also placed on the pressure side near the leading edge. The positions of the uniaxial gauge and the rosettes are illustrated in Figure 7.
Due to blade geometry and the presence of fillets, the gauges could not be placed directly at the stress hotspots, where the strain amplitudes are largest; instead, they were positioned at the nearest locations permitted by the geometrical constraints. The locations of the static and dynamic stress hotspots were previously determined through Computational Fluid Dynamics (CFD) and Finite Element Analysis (FEA) simulations.
In addition, the turbine guide vanes were equipped with a linear encoder to measure their opening angle $o$, while the turbine rotational speed $\omega$ was measured by a proximeter.
@@
The optimization is conducted using strain data measured by one of the branches of rosette 4, located on the pressure side, near the trailing edge and shroud. This sensor measured dynamic strains with the highest potential for optimization, as shown in Figure 8 with strain signals from three different gauges. That potential was determined based on the highest strain cycle during startup, along with the irreducible strain cycle between standstill and the moment when the turbine is ready to be synchronized (speed-no-load).
@@
Other experiments conducted during the measurement campaign allowed us to test $N_{init}=5$ startups for initial trajectories, which included two standard startups and three alternatives designed to cover the space of possible trajectories. The optimization process was run using a budget of $N_{act}=2$ active learning startups and $N_{opt}=1$ optimization trajectory. A last optimization trajectory was run with $T_{st}=60$ s to study the effect of shorter time constraints.
@@
The black-box optimization loop was given a budget of $N_{I}=200$ iterations. NOMAD was launched on 10 parallel threads using the PSD-MADS algorithm (Audet et al., 2008). The various trajectory sampling frequencies were set to balance accurate modeling and time efficiency. Trajectories $\tau_{M}$ were measured at 5,000 Hz and then resampled at $f_{M}=$ 500 Hz. The enveloped trajectories $\tau_{e}$ were calculated with a window $w$ corresponding to 10 s, and then resampled at $f_{e}$ = 10 Hz. On the black-box side, the turbine dynamics simulator created dynamic trajectories $\tau_{D}$ at $f_{D}=10$ Hz, with $f_{S}=f_{D}$ for simulated trajectories $\tau_{S}$.

### v2

@@ 1 Introduction
Among these conditions, HGU startup sequences induce substantial stress on the runner [10]. This poses a significant challenge for hydropower utilities, particularly in light of the anticipated increase in starting cycle frequency [11]. Over the lifetime of the turbine, the stress cycles experienced by the runner contribute to fatigue, leading to crack propagation [12]. The risk of blade cracking increases the frequency of the required inspections, which results in significant financial costs, reduced flexibility, and a loss of power output for hydropower utilities [13]. Therefore, mitigating turbine stresses during startup is crucial to enhance the reliability and longevity of hydroelectric plants in the evolving energy landscape as shown in [14].
Stress levels experienced by hydroturbines during startup are closely linked to the trajectory in terms of guide vane opening and rotational speed as highlighted in both [15] and [16]. The guide vane opening sequence is controlled by the hydraulic turbine governing systems. The governing system is parameterized and ensures that the turbine reaches synchronous speed while respecting specific operational requirements [17]. Optimization of startup sequence parametrization has been addressed by several researchers with different approaches. For example, Seydoux et al. [18] and Muser et al. [19] proposed optimization framework based on Voronoi cell tessellation and deep-learning algorithms, respectively, applied to experimental reduced-scale model data, while Unterluggauer et al. [20] compared the impact of different startup scenario on turbine life expectancy through on-site campaign on a full-scale machine. From a numerical point of view, Schmid et al. [21] proposed an optimization framework combining a runner damage hillchart derived from CFD-FEA simulations with 1D hydraulic transient simulations.
@@
Model testing allows for cost reductions and exploration of a wide variety of startup scenarios without the access and time constraints encountered in prototype tests. However, the transposition of turbine mechanical responses during startup from model to full-scale remains a significant challenge, and no rigorous methodology has been established to date. In contrast, prototype measurements provide insights into the real mechanical behavior of full-scale turbines. However, these measurements are challenging and costly for hydroelectric operators due to time constraints and the harsh environment in which hydroelectric turbines operate, which often leads to the rapid degradation of strain gauges. Therefore, startup optimization must rely on the limited data acquired during short, dedicated measurement campaigns.
@@
Gagnon et al., in 2014 [22] and 2016 [23], presented the results of such campaigns, highlighting the effects of startup sequences on blade stresses. Based on their results, they defined optimal startup scenarios in terms of stress levels. However, these scenarios are mainly derived through trial-and-error processes and expert experience, leaving the potential for further optimization untapped. While optimization may be performed post-campaign using a virtual sensor approach [24], the results cannot be validated without a subsequent experimental campaign.
@@
Given the limited strategies found to optimize startup transients in the literature presented in this study, we propose a novel methodology that allows for the optimization of startup trajectory parameters during a measurement campaign, ultimately contributing to extending the life of hydroturbines. By combining an active learning framework with virtual sensors and black-box optimization, the active learning algorithm sequentially generates various startup parameters to test on the turbine. At each step, new data are used to refine the model and optimization process. This approach was successfully tested during an experimental campaign conducted on a full-scale Francis turbine instrumented with several strain gauge rosettes.
@@
This article is structured as follows. In Section 2, the problem is formulated in details, from HGU dynamics and strain generation to the resulting active optimization problem. Section 3 presents the proposed methodology, which combines an HGU dynamics simulator, virtual strain gauges, black-box optimization and the active learning approach. This method was applied to a real HGU during a measurement campaign, and the results are provided and discussed in Section 4. To facilitate the reading of the paper, a mathematical notation table is provided in appendix A.
@@ 2 Problem description
[…]
@@
On turbine blades, repeated high-amplitude strain cycles lead to fatigue and potential cracks [25]. The greater the cycle amplitude, the greater the accumulated damage to the blade over time. Therefore, in fatigue analysis, the largest amplitude strain/stress cycle, meaning the difference between the highest and lowest strain values over the complete start process, has the most impact on fatigue. As a result, the objective of minimizing fatigue during startup can be approximated by minimizing the amplitude of the largest strain cycle. Formally, this can be expressed as minimizing the following loss over a given trajectory:
$\mathcal{L}(\tau_{M}^{i})=\max_{n\in\nu_{M}^{i}}(s^{i}_{n})-\min_{n\in\nu_{M}^{i}}(s^{i}_{n})$ (4)
@@ 3.1 Startup trajectory simulation
Turbine dynamics can be addressed using the SIMSEN software [26]. A SIMSEN model of the HGU includes the upstream and downstream reservoirs, the water piping system, and the turbine. It can be used to simulate the rotational speed $\omega(t)$ for a predefined guide vane opening $o(t)$ time series. A typical simulation of a single HGU startup takes between 30 seconds and 1 minute. This is problematic for an active optimization framework, where hundreds of simulations must be performed in a limited time. Moreover, in practice, the guide vane opening $o(t)$ is not known beforehand; instead, it is calculated as part of the speed governor model, using the rotational speed $\omega(t)$. To properly address this bidirectional coupling, a co-simulation approach would have to be used in order to ensure communication between the SIMSEN model and speed governor model at each time step, further increasing computational time.
@@ 3.3 Strain modeling with virtual sensors
Evaluating the objective in Equation (16), requires an estimation of $b^{u}$ and $b^{l}$. Given strain measurement data, it is possible to build a model of these distributions using the virtual sensor approach [24]. $b^{u}(\cdot|\omega_{n},o_{n})$ and $b^{l}(\cdot|\omega_{n},o_{n})$ are modeled as conditional normal distributions, parameterized by mean $\mu^{u}_{n}$ and $\sigma^{u}_{n}$ (respectively, $\mu^{l}_{n}$ and $\sigma^{l}_{n}$). The normal distribution was chosen because it is simple and generally a good approximation. This assumption was not validated in the current study. These parameters are estimated using the deep learning model $m_{\phi}(\omega_{n},o_{n})=\left(\hat{\mu}^{u}_{n},\hat{\sigma}^{u}_{n},\hat{\mu}^{l}_{n},\hat{\sigma}^{l}_{n}\right)$ trained on a dataset $\mathcal{D}$ of enveloped trajectories $\{\tau_{e}\}$.
Model $m_{\phi}$ is parameterized as a feed-forward neural network composed of one input layer of dimension 2 ($\omega_{n},o_{n})$, two hidden layers of dimension 32, and one output layer of dimension 4 (two $\mu$ values and two $\sigma$ values). The hidden layers are normalized using batch normalization [27]. As demonstrated in [24], the strain envelope can be effectively predicted given current valve opening and rotational speed values for a given head. Because the trajectories are time series, it could seem intuitive to use recurrent neural networks such as LSTM or GRU and condition the prediction over the past trajectory $\tau_{D}$. However, such architectures do not produce better results experimentally, hinting that, during startups, the current values contain the necessary information to predict the strain envelope.
@@
Since $b^{u}$ and $b^{l}$ are modeled as conditional normal distributions, $m_{\phi}$ was trained using the $\beta$-negative log-likelihood loss function [28], which was shown to robustly estimate heteroscedastic aleatoric uncertainty. Note that predicting $\mu^{u}_{n}$ and $\sigma^{u}_{n}$ is equivalent to estimating the aleatoric uncertainty of $s^{u}_{n}$ (and respectively for $s^{l}_{n}$). Model $m_{\phi}$ was trained over three epochs, using the Adam optimizer with a learning rate of 0.001 and a mini-batch size of 32.
@@ 3.4 Black-box optimization
Given a virtual sensor $m_{\phi}(\omega_{n},o_{n})$, the objective of Equation (16) under the time constraint of Equation (6) can be addressed using black-box optimization. We use the NOMAD 4 [29] with a budget of $N_{I}$ calls to the black-box.
@@
[…]
In the second case, the time constraint is respected. However, $c_{c}$ signals to NOMAD that, given equal strain amplitude, a shorter trajectory is preferred. The last case, representing a violation of the time constraint, is set as an indicator of 1, plus a penalty that increases as the scale of the constraint violation increases to guide NOMAD towards shorter trajectories. If $\omega$ has not stabilized at $\omega_{S}$ after $2\bar{t}_{\mathrm{st}}$, the simulation is stopped and $t^{i}_{\mathrm{st}}$ is set to $2\bar{t}_{\mathrm{st}}$.
To ensure compliance with the time constraint, $c_{c}(\theta^{i})$ in the event of a time constraint violation must be higher than any possible strain-related cost $\alpha_{d}c_{s}(\theta^{i})$. Therefore, $\alpha_{d}$ is set as the inverse of the difference between the maximum and minimum strain values encountered in the whole dataset $\mathcal{D}$. This ensures that $\alpha_{d}c_{s}(\theta^{i})\leq 1$ in all cases while $c_{c}(\theta^{i})\geq 1$ if $t^{i}_{\mathrm{st}}\geq\bar{t}_{\mathrm{st}}$. The different components of the black box are illustrated in Figure 6.
Figure 6: Different components of the black box given measured trajectories dataset $\mathcal{D}$
@@ 3.5 Active learning and optimization
The proposed active scheme is presented in Figure 7, where the black box optimizer is used at every outer loop iteration to provide the parameters $\theta^{j+1}$ to test on the turbine. The new measured trajectory is then added to the dataset, and the virtual sensor is retrained with the new data. The process goes through three phases with predetermined trajectory budgets: initialization ($N_{init}$), active learning ($N_{act}$), and optimization ($N_{opt}$), such that $N_{init}+N_{act}+N_{opt}=N_{\mathrm{b}}$. For the remainder of the paper, the variables indexed by $i$ in the single black-box loop will be indexed by i,j{i,j} to include nested loop index $j$.
@@
The objective of the active learning phase is to explore the space of trajectories in order to supply the most relevant data for training the virtual sensor. To do so, the same black-box optimization system as presented in Section 3.4 is used. However, the evaluator is modified to account for the epistemic uncertainty, expressed as the standard deviation $\sigma_{\mathrm{ep},n}^{\{u,l\},i,j}$ of the ensemble’s individual $m$ model predictions of sum $\hat{\mu}^{\{u,l\},i,j}_{n}+\hat{\sigma}^{\{u,l\},i,j}_{n}$. Inspired by the upper confidence bound approach [30] for exploration, the strain cost in Equation (18) is modified as follows:
@@ 4.1 Experimental setup
Our case study is an instrumented Francis turbine runner installed in a storage hydroelectric power plant. The specific speed of the turbine is $n_{QE}=nQ^{0.5}/(gH)^{3/4}=65$, with $n$, the runner frequency, $Q$, the nominal flow rate, $g$, the gravity and $H=72$ m, the nominal head. For this power plant, the time limit for reaching synchronous speed was set to $\bar{t}_{\mathrm{st}}=90$s.
Two blades were instrumented with four strain gauge rosettes near the trailing edge, with two rosettes positioned on the suction side and two on the pressure side. An additional uniaxial strain gauge was also placed on the pressure side near the leading edge. The positions of the uniaxial gauge and the rosettes are illustrated in Figure 8. The typical setup for data acquisition for such measurement campaign, where sensor data on the rotating part can be transmitted either via slip-ring or through a Wi-Fi telemetry system for on-site processing, can be found in [31].
Due to blade geometry and the presence of fillets, the gauges could not be placed directly at the stress hotspots, where the strain amplitudes are largest; instead, they were positioned at the nearest locations permitted by the geometrical constraints. The identification of stress hotspots location was based on CFD and FEA simulations using state-of-the-art numerical procedures as described in [32]. In the turbine used as test-case, two static stress hotspots were identified near the trailing edge of the blades, at the fillets between the shroud / hub and the blades, which are typical hotspot position for Francis turbines.
The turbine guide vanes were equipped with a linear encoder to measure their opening angle $o$, while the turbine rotational speed $\omega$ was measured by a proximeter.
@@
The optimization is conducted using strain data measured by gauge 2 of rosette 4, located on the pressure side, near the trailing edge and shroud. This sensor measured dynamic strains with the highest potential for optimization, as shown in Figure 9 with strain signals from three different gauges. That potential was determined based on the highest strain cycle during startup, along with the irreducible strain cycle between standstill and the moment when the turbine is ready to be synchronized (speed-no-load).
@@
Other experiments conducted during the measurement campaign allowed us to test $N_{init}=5$ startups for initial trajectories, which included two standard startups and three alternatives designed to cover the space of possible trajectories. The optimization process was run using a budget of $N_{act}=2$ active learning startups and $N_{opt}=1$ optimization trajectory. A last optimization trajectory was run with $\bar{t}_{\mathrm{st}}=60$ s to study the effect of shorter time constraints.
@@
The black-box optimization loop was given a budget of $N_{I}=200$ iterations. NOMAD was launched on 10 parallel threads using the PSD-MADS algorithm [33]. The various trajectory sampling frequencies were set to balance accurate modeling and time efficiency. Trajectories $\tau_{M}$ were measured at 5,000 Hz and then resampled at $f_{M}=$ 500 Hz. The enveloped trajectories $\tau_{e}$ were calculated with a window $w$ corresponding to 10 s, and then resampled at $f_{e}$ = 10 Hz. On the black-box side, the turbine dynamics simulator created dynamic trajectories $\tau_{D}$ at $f_{D}=10$ Hz, with $f_{S}=f_{D}$ for simulated trajectories $\tau_{S}$.